OpenStudy (anonymous):

x dy/dx-y=sinx〖.y〗^2

ganeshie8 (ganeshie8):

divide x through out, its a bernouli equation too familiar with solving these ?

ganeshie8 (ganeshie8):

\[\rm x\dfrac{dy}{dx} - y = (\sin x )(y^2)\] like this ?

ganeshie8 (ganeshie8):

dividign `xy^2` through out, you get : \[\rm \frac{1}{y^2}\dfrac{dy}{dx} -\frac{1}{x}\frac{1}{y} = \frac{\sin x}{x} \]

ganeshie8 (ganeshie8):

substitute \(\large \rm \frac{1}{y} = v \implies \frac{1}{y^2}\frac{dy}{dx} = -\frac{dv}{dx}\)

ganeshie8 (ganeshie8):

the equation becomes : \[\rm -\frac{dv}{dx} - \dfrac{1}{x}v = \dfrac{\sin x}{x}\]

ganeshie8 (ganeshie8):

which is same as : \[\rm \frac{dv}{dx} + \dfrac{1}{x}v = -\dfrac{\sin x}{x}\]

ganeshie8 (ganeshie8):

its a linear equation now, find the integrating factor and see if u can take it home

OpenStudy (anonymous):

k i got it thanq very much

ganeshie8 (ganeshie8):

yw:)